OpenStudy (amy0799):

Find equations of the two tangent lines to the graph of f(x) = 2x^2 that pass through the point (0, -2). line with positive slope: line with negative slope:

OpenStudy (michele_laino):

the equation of the tangent line, with slope m and which passes at the point (0,-2), is: \[\Large y - \left( { - 2} \right) = m\left( {x - 0} \right)\]

OpenStudy (amy0799):

I understand that, but how do I get the positive and negative slope?

OpenStudy (michele_laino):

I simplify that equation like below: \[\Large y = mx - 2\] then I consider this algebraic system: \[\Large \left\{ \begin{gathered} y = mx - 2 \hfill \\ y = 2{x^2} \hfill \\ \end{gathered} \right.\] I solve it, substitution method, namely I replace y into the second equation, by mx-2, so I get this: \[\Large mx - 2 = 2{x^2}\] now, in order to get the requested values for m, we have to request that the discriminant of that equation is equal to zero

OpenStudy (amy0799):

i found the derivative to be 4x. and plugging in 0 to 4x gets me o. so would the equations be: y+2=0(x-0) y+2=-0(x-0)

OpenStudy (michele_laino):

I rewrite my last equation as below: \[\Large 2{x^2} - mx + 2 = 0\] which is a quadratic equation. Now I request that its discriminant, \Delta is equal to zero. \[\Large \Delta = {b^2} - 4ac = {\left( { - m} \right)^2} - 4 \cdot 2 \cdot 2 = 0\] Please, solve for m

OpenStudy (amy0799):

is m=-4?

OpenStudy (michele_laino):

yes! nevertheless it is one solution, we have to find the second solution

OpenStudy (michele_laino):

\[\Large {m^2} - 16 = 0\] what is m?

OpenStudy (amy0799):

4

OpenStudy (michele_laino):

correct! your tangent lines are: y=4x-2 y=-4x-2

OpenStudy (amy0799):

thank you so much!

OpenStudy (michele_laino):

:)